OpenStudy (anonymous):

Find the equation of the line that passes through the point (2,11) and is parallel to y=4x+4

jimthompson5910 (jim_thompson5910):

The equation " is parallel to y=4x+4" so this equation has a slope of 4 (parallel lines have equal slopes)

jimthompson5910 (jim_thompson5910):

we know m = 4

jimthompson5910 (jim_thompson5910):

" the line that passes through the point (2,11)" so x = 2 y = 11

jimthompson5910 (jim_thompson5910):

we have this information m = 4 x = 2 y = 11 Plug that into y = mx+b and solve for b

jimthompson5910 (jim_thompson5910):

I'll let you do that

OpenStudy (anonymous):

@jim_thompson5910 I am in a hurry please explain more/

jimthompson5910 (jim_thompson5910):

it's not good to be rushed, so go do what you need to do and come back

OpenStudy (anonymous):

formula is y-y=m(x-x) the answer is y=4x+3